# Probing soft signals of gravitational-wave memory with space-based interferometers

### 30 Mar 2026 v1

@@ III.4 Constraint on the source from the displacement-memory amplitude
From Eq. (33), the total offset of vacuum null displacement memory at future null infinity in an asymptotically flat spacetime can be expressed by $\Delta h=\frac{2\pi}{R}\bar{\eth}^{2}\mathfrak{D}^{-1}\frac{dE_{\text{gw}}}{d\Omega_{2}}$, where $\bar{\eth}$ is the spin-lowering operator, $\mathfrak{D}=D^{2}(D^{2}+2)/8$, and $D^{2}$ is the spherical Laplacian. Including the ordinary contribution, the result is [67]
$\Delta h=\frac{1}{2}\bar{\eth}^{2}\mathfrak{D}^{-1}\Delta Q,$ (34)
with $\Delta Q=\left(\Delta m+4\pi\frac{dE_{\text{gw}}}{d\Omega_{2}}\right)\frac{1}{R}$, where $\Delta m$ is the change of Bondi mass aspect. For $l\geq 2$, Eq. (34) reads
$(\Delta h)_{lm}=4\sqrt{\frac{(l-2)!}{(l+2)!}}(\Delta Q)_{lm}.$ (35)
where $J_{lm}$ denotes the expansion coefficient of a spin-$s$ field $J(\Theta,\Phi)$ in terms of $_{s}Y_{lm}(\Theta,\Phi)$. Here we note that a constraint on the source can be derived from $|\Delta h|$,
$\mathcal{Q}$ $\equiv\int d\Omega_{2}\,|\Delta Q|^{2}-\sum_{l=0,1}\sum_{m}|(\Delta Q)_{lm}|^{2}$ (36)
$=\sum_{l\geq 2}\sum_{m}|(\Delta Q)_{lm}|^{2}\geq\frac{|\Delta h|^{2}}{\mathcal{I}(\Theta,\Phi)},$
with $\mathcal{I}=\sum_{l\geq 2}^{\infty}\sum_{m}\frac{16(l-2)!}{(l+2)!}\left|{}_{-2}Y_{lm}\right|^{2}$. For $\Theta=0$, $_{-2}Y_{lm}(\Theta,\Phi)=e^{im\Phi}\sqrt{\frac{2l+1}{4\pi}}\delta_{m,2}$, $\mathcal{I}=4/(3\pi)$, thus $\mathcal{Q}\geq(3\pi/4)|\Delta h|^{2}$. Taking into account the cosmic expansion discussed in Sec. III.1, this bound becomes $\mathcal{Q}_{R\to d_{L}}\geq(3\pi/4)|\Delta h|^{2}/(1+z)^{2}$.
@@ IV.1 TDI response
The propagation of laser light between the SC can be described in the leading-order eikonal approximation, as presented in Sec. III.1. The light frequency $\omega_{s}$ at $\mathbf{x}_{s}(t_{s})$ is fixed by the laser source, while the light frequency $\omega_{r}$ measured locally at $\mathbf{x}_{r}(t)$ is affected by both the motion of $\mathbf{x}_{r}$ [see Eq. (79) in Appendix B] and the light propagation. The effect of GW on the one-way Doppler shift $y_{rs}(t)\equiv[\omega_{r}(t)-\omega_{s}]/\omega_{s}$ can be analyzed conveniently in the TT coordinates, with the relevant metric given by Eq. (1). For light propagating between two static test-masses, the relative frequency modulation comes solely from the GW-induced modification of the photon dispersion relation, which for a plane wave $h_{ij}(t,\mathbf{x})=h_{ij}(u=t-\hat{\mathbf{k}}\cdot\mathbf{x})$ is given by Eq. (20):
@@
where the time delay operators are defined by
$D_{ij}x(t)$ $\equiv x(t-L_{ij}),$ (39)
$D_{i_{1}\cdots i_{n}}x(t)$ $\equiv\prod_{k=1}^{n-1}D_{i_{k}i_{k+1}}x(t),$ (40)
and for the computation of GW signals, it suffices to make the following approximation:
$D_{i_{1}\cdots i_{n}}x(t)\approx x\left(t-\sum_{k=1}^{n-1}L_{i_{k}i_{k+1}}\right).$ (41)
@@
The TDI response is further simplified in the static equal-armlengths approximation (SEA), in which we neglect the motion of SC and use $L_{rs}\approx L=\text{const}$. The static-detector approximation is reasonable for a short-term observation, and the equal-armlengths approximation can also be sufficiently accurate in such cases, since the armlengths vary only at the percent level. Given the time-domain waveform $h_{ij}(t)=h_{ij}(t,\mathbf{0})$, the frequency spectrum of the time-domain TDI signals (with a sufficiently high sampling frequency) can be easily derived in this approximation. Using Eq. (37), the result for (38) is
$\tilde{X}_{1}(f)$ $=\frac{1}{2}\left[n_{13}^{i}n_{13}^{j}\left(F_{13}+e^{2\pi ifL}F_{31}\right)-(3\leftrightarrow 2)\right]\tilde{h}_{ij},$ (42)
where
@@
$\tilde{X}_{1}(f)\approx 8\pi^{2}L^{2}f^{2}\sum_{\lambda}\left(n_{13}^{i}n_{13}^{j}-n_{21}^{i}n_{21}^{j}\right)\tilde{h}_{ij}\,e^{2\pi if\,\hat{\mathbf{k}}\cdot\mathbf{X}},$ (44)
@@
For the long-term observation using a single detector or the joint observation using multiple detectors, we parameterize $\{\mathbf{a,b,\hat{k}}\}$ in the solar-system-barycentric (SSB) ecliptic coordinate frame, as depicted in the right panel of Fig. 8. In the equal-armlengths approximation, the positions of SC in the SSB frame can be obtained as
$x_{Ii}^{\prime}=X^{\prime}_{i}(\eta)+R_{ij}(\eta)\,x_{Ij},$ (47)
with the guiding center $\mathbf{X}^{\prime}=R(\cos\eta,\sin\eta,0)$ and $R_{ij}$ a rotation matrix that depends on a phase constant $\Phi_{0}$ (corresponding to $\eta_{0}+\xi_{0}$ in [61]):
@@ IV.2 Soft memory signals
Figure 9: Comparison between analytical and numerical spectrum of $A_{1}(t)$ for the arctan model (7) at LISA, with parameters $H=10^{-19}$, $T=0.05\,\text{yr}$, $t_{*}=0.5T$, $\{\beta,\lambda,\psi\}=\{-1.4,2.3,\pi/3\}$, and $f_{*}\in\{10^{-5},0.01,1,5\}\,\text{Hz}$. The analytical spectrum (black dashed line) is obtained in the SEA, with the soft limit indicated by the blue dashed line. The numeric spectrum (colored solid line) is computed using fastlisaresponse with the equal-armlength orbit and a sampling frequency of $1\,\text{Hz}$. The vertical dashed lines correspond to integer multiples of $f=1/(2L)$. The time-domain waveforms are shown in the figure inset.
In the SEA, the frequency-domain signals $\{\tilde{A}_{1},\tilde{E}_{1},\tilde{T}_{1}\}$ can be calculated using Eqs. (42) and (45). The signals are degenerate under the transformation from $\{\theta,\psi\}$ to $\{\theta^{\prime},\psi^{\prime}\}=\{\pi-\theta,\pi-\psi\}$, and merely changes sign under the transformation from $\psi$ to $\psi^{\prime}=\psi\pm\pi/2$. In the low-frequency approximation (44), there is an additional degeneracy under $\phi\leftrightarrow\phi+\pi$. Moreover, the signals in the SEA at $f\to 0$ depend only on four parameters, while a soft memory signal has five parameters in total. The measurement of a soft memory event with a single detector is therefore possible precisely because the response function deviates from its low-frequency approximation. When this deviation is small, the signal suffers from a stronger degeneracy.
@@ IV.3 Optimal signal-to-noise ratio
The logarithmic spectral density of soft displacement-memory and velocity-memory signal for $\phi=\theta=\psi=0$ are shown in Figs. 10 and 11, respectively. The spectral density is normalized such that the peak value equals unity. We also show results obtained using the low-frequency approximation of the response function as dashed lines; these correspond to the flat signal spectrum at the low-frequency end in the bottom panel. As can be seen, a significant portion of the spectrum cannot be described using the low-frequency approximation in the case of displacement memory. Meanwhile, the low-frequency approximation is much better for Taiji in the case of velocity memory, indicating a strong degeneracy of the signal.
@@ IV.4 Distinguishability of real linear and quadratic corrections to the soft displacement-memory signal
As discussed in Sec. II.4, there is necessarily a correction to the soft waveform of displacement memory in the form of Eq. (15). However, this cannot be faithfully distinguished unless the SNR is sufficiently large. As concrete examples, in this section we estimate the distinguishability of real linear and quadratic corrections, corresponding to $C(f)=a_{0}f$ and $C(f)=a_{1}f^{2}$, respectively, with $a_{0,1}\in\mathbb{R}$. In the following, we consider the cases of linear and quadratic corrections separately.
@@
where $D$ is the number of parameters in the waveform template. In the present case, we choose $h_{1}$ ($h_{2}$) to be the waveform with (without) the correction, and take $D=6$. Note that for a generic correction factor $C(f)\in\mathbb{C}$, $\mathcal{M}$ depends only on the integrals $\sum_{U}\int d\ln f\left[\frac{4f|\tilde{U}_{a}|^{2}}{S_{U}(f)}\right]\text{Re}\,C$ and $\sum_{U}\int d\ln f\left[\frac{4f|\tilde{U}_{a}|^{2}}{S_{U}(f)}\right]|C|^{2}$. Fig. 14 shows the estimated critical values of $a_{0,1}$ distinguishable by Taiji and BBO as functions of the SNR, which are insensitive to $\{\phi,\theta,\psi\}$. In the next section, we will examine the impact of such corrections to the parameter estimation using the uncorrected soft-waveform template in concrete examples.
@@ V Parameter estimation simulation
However, the FIM estimate cannot be reliably trusted at low SNR, which is likely a realistic regime. In the following, we present results from Bayesian parameter-estimation simulations, focusing on the case of displacement memory. For sampling the posterior, we use the MCMC sampler emcee [34] and plot the posterior distributions using corner [35]. For concreteness, we consider the measurements using Taiji, LISA-Taiji network (described in Fig. 8) and BBO. For LISA-Taiji, we set $\eta_{\text{Taiji}}=\eta_{\text{LISA}}+2\pi/9$, $\Phi_{0}=0$, and $\eta_{\text{LISA}}=\pi/4$; the log-likelihood is taken as the sum of those of LISA and Taiji. We use the parameter set $\boldsymbol{\xi}=\{H,\tau_{*},\lambda,\sin\beta,\Psi\}$ for LISA-Taiji, and $\boldsymbol{\xi}=\{H,\tau_{*},\phi,\cos\theta,\psi\}$ for a single BBO detector. The priors are taken to be uniform, as listed in Table 3.
@@
Figure 21: Spectra of $2\pi f|\tilde{h}(f)|/H=1+C(f)$ in various scenarios. The solid and dashed lines represent the real linear and quadratic corrections considered in Fig. 19, respectively. The blue lines correspond to $|a_{0}|=50\,\text{Hz}^{-1}$ or $|a_{1}|=20000\,\text{Hz}^{-2}$, the red lines correspond to $|a_{0}|=10\,\text{Hz}^{-1}$ or $|a_{1}|=10000\,\text{Hz}^{-2}$. The gray and black dot-dashed lines correspond to the waveforms of compact binary scattering and merger events considered in Fig. 20, respectively.
@@
Fig. 16 shows a simulated measurement result by LISA-Taiji in the high-SNR regime, with $\hat{H}=10^{-19}$. As can be seen, $\tau_{*}$ is measured with significantly higher precision (this is partly because the considered signal duration is relatively long), while the memory amplitude $H$ is better constrained than the angular parameters $\{\lambda,\beta,\Psi\}$. Joint observations with LISA and Taiji is found to greatly improve the measurement accuracy compared with observations using Taiji alone, even though the SNR is similar. We also test the stochasticity of the noise injection and find that, while it does not significantly affect the results, it introduces an overall bias. In the absence of noise injection, the results agree well with the FIM estimation.
@@
The lower-SNR regime is examined in Fig. 17. We find that even for $\rho\sim 10$, observations with Taiji alone can still effectively constrain the parameters, in particular $H$ and $\tau_{*}$. Notably, the FIM substantially overestimates the measurement error when the SNR is low. If the sky location of the source is known a priori (from other observations), the measurement precision improves dramatically. For the soft velocity-memory signal with the same SNR, the measurement errors are slightly larger. The impact of source direction on the measurement precision is examined in the upper panel of Fig. 17, where the source colatitude $\theta$ in the detector frame of Taiji is varied while the other parameters and the SNRs are kept fixed at 13.7. The measurement errors are found to be of the same order of magnitude, although $\sigma_{s_{\beta}}$ is relatively large for $\hat{\theta}=0.01\pi$, likely due to the signal degeneracy at $\theta=0$.
@@
The bottom panel of Fig. 18 shows the simulated measurement results by a single BBO detector at different SNRs. We again find that the measurement precision of $H$ and $\tau_{*}$ is better when the SNR is sufficiently high. Compared with measurements using Taiji, the results agree better with the FIM estimates for $\rho\sim 10$, possibly due to weaker parameter degeneracies of the signal at BBO.
@@
Figs. 19-20 present example results of parameter estimation for corrected soft waveforms using an uncorrected template. In Fig. 19, we show the relative measurement errors of a soft displacement-memory signal with real linear and quadratic corrections (as discussed in Sec. IV.4), i.e., $\tilde{h}=\frac{iHe^{-2i\psi}}{2\pi f}e^{2\pi ift_{*}}[1+C(f)]$ with $C(f)=a_{0}f+a_{1}f^{2}$. To isolate the intrinsic parameter-estimation performance, we omit noise injection. The considered correction factors are relatively large and are plotted in Fig. 21. We find that these corrections only have a minor impact on the measurement of $\tau_{*}$, but they introduce clear biases in $\log_{10}H$. When the correction is moderate, however, the bias remains modest, staying within the $2\sigma$ deviation. For the other parameters, which have considerably lower measurement precision and show visible bias even in the absence of corrections, the results are not significantly affected, although the biases tend to increase.
@@
As realistic examples, we simulate the measurement of a stellar-mass compact binary scattering event with the Taiji detector and a quasi-circular stellar-mass BBH merger event with the BBO detector. For the former, we use the waveform in Fig. 4, and for the latter, the $q=1$ non-spinning waveform in Fig. 7 (taking into account the comic expansion). As discussed previously, both waveforms can be approximated by soft waveforms of displacement memory with real correction factors $C(f)$, shown in Fig. 21. As above, a soft-waveform template is employed to perform the parameter estimation. The posterior distributions of the memory amplitude and polarization angle, together with the detailed parameter settings, are shown in Fig. 20. In both examples, we find that the memory parameters can be effectively constrained, with the amplitude $H$ measured with a higher precision.

### 13 Apr 2026 v2

@@ III.4 Constraint on the source from the displacement-memory amplitude
From Eq. (27), the total offset of vacuum null displacement memory at future null infinity in an asymptotically flat spacetime can be expressed by $\Delta h=\frac{2\pi}{R}\bar{\eth}^{2}\mathfrak{D}^{-1}\frac{dE_{\text{gw}}}{d\Omega_{2}}$, where $\bar{\eth}$ is the spin-lowering operator, $\mathfrak{D}=D^{2}(D^{2}+2)/8$, and $D^{2}$ is the spherical Laplacian. Including the ordinary contribution, the result is [72] $\Delta h=\frac{1}{2}\bar{\eth}^{2}\mathfrak{D}^{-1}\Delta Q$, with $\Delta Q=\left(\Delta m+4\pi\frac{dE_{\text{gw}}}{d\Omega_{2}}\right)\frac{1}{R}$, where $\Delta m$ is the change of Bondi mass aspect. [^10] E.g., for massive particles at timelike infinity, the Bondi mass aspect may be taken as $m(\Theta,\Phi)=\sum_{I}\frac{(1-v_{I}^{2})^{3/2}}{[1-\mathbf{v}_{I}\cdot\mathbf{n}(\Theta,\Phi)]^{3}}m_{I}$ [78]. Substituting this in Eq. (28) gives Eq. (24). For $l\geq 2$, this gives
$(\Delta h)_{l,m}=4\sqrt{\frac{(l-2)!}{(l+2)!}}(\Delta Q)_{l,m}.$ (28)
where $J_{l,m}$ denotes the expansion coefficient of a spin-$s$ field $J(\Theta,\Phi)$ in terms of $_{s}Y_{lm}(\Theta,\Phi)$. Here we note that a constraint on the source can be derived from $|\Delta h|$,
$\mathcal{Q}$ $\equiv\int d\Omega_{2}\,|\Delta Q|^{2}-\sum_{l=0,1}\sum_{m}|(\Delta Q)_{l,m}|^{2}$ (29)
$=\sum_{l\geq 2}\sum_{m}|(\Delta Q)_{l,m}|^{2}\geq\frac{|\Delta h|^{2}}{\mathcal{I}(\Theta,\Phi)},$
with $\mathcal{I}=\sum_{l\geq 2}^{\infty}\sum_{m}16\frac{(l-2)!}{(l+2)!}\left|{}_{-2}Y_{lm}\right|^{2}$. Since $\sum_{m}|_{-s}Y_{lm}|^{2}=(2l+1)/(4\pi)$, $\mathcal{I}=4/(3\pi)$, thus $\mathcal{Q}\geq(3\pi/4)|\Delta h|^{2}$. Taking into account the cosmic expansion (see Sec. III.1), this bound becomes $\mathcal{Q}_{R\to d_{L}}\geq(3\pi/4)|\Delta h|^{2}/(1+z)^{2}$.
@@ IV.1 TDI response
The propagation of laser light between the SC can be described in the leading-order eikonal approximation, as presented in Sec. III.1. The light frequency $\omega_{s}$ at $\mathbf{x}_{s}(t_{s})$ is fixed by the laser source, while the light frequency $\omega_{r}$ measured locally at $\mathbf{x}_{r}(t)$ is affected by both the motion of $\mathbf{x}_{r,s}$ [see Eq. (64) in Appendix B] and the light propagation. The effect of GW on the one-way Doppler shift $y_{rs}(t)\equiv[\omega_{r}(t)-\omega_{s}]/\omega_{s}$ can be analyzed conveniently in the TT coordinates, with the relevant metric given by Eq. (1). For light propagating between two static test-masses, the relative frequency modulation comes solely from the GW-induced modification of the photon dispersion relation, which for a plane wave $h_{ij}(t,\mathbf{x})=h_{ij}(u=t-\hat{\mathbf{k}}\cdot\mathbf{x})$ is given by Eq. (18):
@@
where the time delay operators are defined by $D_{ij}x(t)\equiv x(t-L_{ij})$ and $D_{i_{1}\cdots i_{n}}x(t)\equiv\prod_{k=1}^{n-1}D_{i_{k}i_{k+1}}x(t)$. For the computation of GW signals, it suffices to make the following approximation: $D_{i_{1}\cdots i_{n}}x(t)\approx x\left(t-\sum_{k=1}^{n-1}L_{i_{k}i_{k+1}}\right)$.
@@
The TDI response is further simplified in the static equal-armlengths approximation (SEA), in which we neglect the motion of SC and use $L_{rs}\approx L=\text{const}$. The static-detector approximation is reasonable for a short-term observation, and the equal-armlengths approximation can also be sufficiently accurate in such cases, since the armlengths vary only at the percent level. Given the time-domain waveform $h_{ij}(t)=h_{ij}(t,\mathbf{0})$, the frequency spectrum of the time-domain TDI signals (with a sufficiently high sampling frequency) can be derived in this approximation by using Eq. (30). The result for (31) is
$\tilde{X}_{1}$ $=\frac{1}{2}\left[n_{13}^{i}n_{13}^{j}\left(F_{13}+e^{2\pi ifL}F_{31}\right)-(3\leftrightarrow 2)\right]\tilde{h}_{ij},$ (32)
with $\tilde{h}_{ij}(f)=\frac{1}{2}(e_{ij}^{+}+ie_{ij}^{\times})\tilde{h}(f)+\frac{1}{2}(e_{ij}^{+}-ie_{ij}^{\times})[\tilde{h}(-f)]^{*}$ ($\tilde{h}(-f)]^{*}$ is the Fourier transform of $h^{*}(t)$), and
@@
$\tilde{X}_{1}\approx 8(\pi fL)^{2}\sum_{\lambda}\left(n_{13}^{i}n_{13}^{j}-n_{21}^{i}n_{21}^{j}\right)\tilde{h}_{ij}\,e^{2\pi if\,\hat{\mathbf{k}}\cdot\mathbf{X}},$ (34)
@@
For the long-term observation using a single detector or the joint observation using multiple detectors, we parameterize $\{\mathbf{a,b,\hat{k}}\}$ in the solar-system-barycentric (SSB) ecliptic coordinate frame, as depicted in the right panel of Fig. 8. In the equal-armlengths approximation, the positions of SC in the SSB frame can be obtained as $x_{Ii}^{\prime}=X^{\prime}_{i}(\eta)+R_{ij}(\eta)\,x_{Ij}$, with the guiding center $\mathbf{X}^{\prime}=R(\cos\eta,\sin\eta,0)$ and $R_{ij}$ a rotation matrix that depends on a phase constant $\Phi_{0}$ (corresponding to $\eta_{0}+\xi_{0}$ in [66]):
@@ IV.2 Soft memory signals
Figure 9: Comparison between analytical and numerical spectrum of $A_{1}(t)$ for the arctan model (7) at LISA, with parameters $H=10^{-19}$, $T=0.05\,\text{yr}$, $t_{*}=0.5T$, $\{\beta,\lambda,\psi\}=\{-1.4,2.3,\pi/3\}$, and $f_{*}\in\{10^{-5},0.01,1,5\}\,\text{Hz}$. The analytical spectrum (black dashed line) is obtained in the SEA, with the soft limit indicated by the blue dashed line. The numerical spectrum (colored solid line) is computed using fastlisaresponse with the equal-armlength orbit and a sampling frequency of $1\,\text{Hz}$. The vertical dashed lines correspond to integer multiples of $f=1/(2L)$. The time-domain waveforms are shown in the figure inset.
In the SEA, the frequency-domain signals $\{\tilde{A}_{1},\tilde{E}_{1},\tilde{T}_{1}\}$ can be calculated using Eqs. (32) and (35). The signals are degenerate under the transformation from $\{\theta,\psi\}$ to $\{\theta^{\prime},\psi^{\prime}\}=\{\pi-\theta,\pi-\psi\}$, and merely changes sign under the transformation from $\psi$ to $\psi^{\prime}=\psi\pm\pi/2$. In the low-frequency approximation (34), there is an additional degeneracy under $\phi\leftrightarrow\phi+\pi$. Moreover, the signals in the SEA at $f\to 0$ depend only on four parameters, while a soft memory signal has five parameters in total. The measurement of a soft memory event with a single detector is therefore possible precisely because the response function deviates from its low-frequency approximation. When this deviation is small, the signal suffers from a strong degeneracy.
@@ IV.3 Optimal signal-to-noise ratio
The logarithmic spectral density distributions of the squared SNR for the soft displacement-memory and velocity-memory signals with $\phi=\theta=\psi=0$ are shown in Figs. 10 and 11, respectively (the squared SNR is proportional to the area under each curve). The spectral density is normalized such that the peak value equals unity. We also show results obtained using the low-frequency approximation of the response function as dashed lines; these correspond to the flat signal spectrum at the low-frequency end in the bottom panel. As can be seen, a significant portion of the spectrum cannot be described using the low-frequency approximation in the case of displacement memory. Meanwhile, the low-frequency approximation is much better for Taiji in the case of velocity memory, indicating a strong degeneracy of the signal.
@@ IV.4 Distinguishability of real linear and quadratic corrections to the soft displacement-memory signal
As discussed in Sec. II.4, there is necessarily a correction to the soft waveform of displacement memory in the form of $\tilde{h}(f)=\frac{iHe^{-2i\psi}}{2\pi f}e^{2\pi ift_{*}}[1+C(f)]$. However, this cannot be faithfully distinguished unless the SNR is sufficiently large. As concrete examples, in this section we estimate the distinguishability of real linear and quadratic corrections, corresponding to $C(f)=a_{0}|f|$ and $C(f)=a_{1}f^{2}$, respectively, with $a_{0,1}\in\mathbb{R}$.
@@
where $D$ is the number of parameters in the waveform template. In the present case, we choose $h_{1}$ ($h_{2}$) to be the waveform with (without) the correction, and take $D=6$. Note that if $C(f)=[C(-f)]^{*}$, then $\tilde{h}_{ij}^{\text{corrected}}=\tilde{h}_{ij}^{\text{uncorrected}}[1+C(f)]$. Fig. 14 shows the estimated critical values of $a_{0,1}$ distinguishable by Taiji and BBO as functions of the SNR, which are insensitive to $\{\phi,\theta,\psi\}$. In the next section, we will examine the impact of such corrections to the parameter estimation using the uncorrected soft-waveform template in concrete examples.
@@ V Parameter estimation simulation
However, the FIM estimate cannot be reliably trusted at low SNR, which is likely a realistic regime. In the following, we present results from Bayesian parameter-estimation simulations, focusing on the case of displacement memory. For sampling the posterior, we use the MCMC sampler emcee [38] and plot the posterior distributions using corner [39]. For concreteness, we consider the measurements using Taiji, LISA-Taiji network (described in Fig. 8) and BBO. For LISA-Taiji, we set $\eta_{\text{Taiji}}=\eta_{\text{LISA}}+2\pi/9$, $\Phi_{0}=0$, and $\eta_{\text{LISA}}=\pi/4$; the log-likelihood is taken as the sum of those of LISA and Taiji. As mentioned in Sec. IV.1, the signals at a single detector are degenerate under the transformation $\{\theta,\psi\}\leftrightarrow\{\pi-\theta,\pi-\psi\}$; this degeneracy can be removed by restricting $\psi\in[0,\pi/2]$; it cannot be easily removed in the SSB-frame parameterization. We therefore adopt the parameter set $\boldsymbol{\xi}=\{H,\tau_{*},\lambda,\sin\beta,\Psi\}$ for LISA-Taiji, and $\boldsymbol{\xi}=\{H,\tau_{*},\phi,\cos\theta,\psi\}$ for a single detector. The priors are taken to be uniform, as listed in Table 3.
@@
Figure 21: Spectra of $2\pi f|\tilde{h}(f)|/H=1+C(f)$ in various scenarios. The solid and dashed lines represent the real linear and quadratic corrections considered in Fig. 19, respectively. The blue lines correspond to $|a_{0}|=50\,\text{Hz}^{-1}$ or $|a_{1}|=20000\,\text{Hz}^{-2}$, the red lines correspond to $|a_{0}|=10\,\text{Hz}^{-1}$ or $|a_{1}|=10000\,\text{Hz}^{-2}$. The gray and black lines correspond to the waveforms of compact binary scattering and merger events considered in Fig. 20, respectively. The gray dot-dashed line shows $1+C(f)$, while the gray dashed line shows 1+C(f)1+C(-f); in the other cases, C(f)=C(f)C(f)=C(-f).
@@
Fig. 16 shows a simulated measurement result by LISA-Taiji in the high-SNR regime, with $\hat{H}=10^{-19}$. As can be seen, $\tau_{*}$ is measured with significantly higher precision (this is partly because the considered signal duration is relatively long), while the memory amplitude $\log_{10}H$ is better constrained than the angular parameters $\{\lambda,\sin\beta,\Psi\}$. Joint observations with LISA and Taiji is found to greatly improve the measurement accuracy compared with observations using Taiji alone, even though the SNR is similar. We also test the stochasticity of the noise injection and find that, while it does not significantly affect the results, it introduces an overall bias. In the absence of noise injection, the results agree well with the FIM estimation.
@@
The lower-SNR regime is examined in Fig. 17. We find that even for $\rho\sim 10$, observations with Taiji alone can still effectively constrain the parameters, in particular $\log_{10}H$ and $\tau_{*}$. If the sky location of the source is known a priori from other observations, the measurement precision improves dramatically. For the soft velocity-memory signal with the same SNR, the measurement errors are slightly larger. Joint observations with LISA–Taiji can again greatly improve parameter estimation; a concrete example is presented in Table 4.
$\log_{10}H$ $\lambda$ $\sin\beta$ $\Psi$ $\tau_{*}$
20-20 4.5 0.51 0.12 0.7
$19.98^{+0.03}_{-0.03}$ $4.51^{+0.05}_{-0.03}$ $0.58^{+0.16}_{-0.19}$ $0.15^{+0.04}_{-0.04}$ $0.701^{+0.002}_{-0.002}$
0.027 0.035 0.21 0.037 0.0019
Table 4: Example parameter-estimation results for a soft displacement memory signal with LISA–Taiji. The first row lists the true values, while the last row gives the FIM estimates of the measurement errors. The optimal SNR is 16.
The dependence of measurement precision on the angular parameters for a single detector is illustrated in the upper panel of Fig. 17, where the source colatitude $\theta$ and polarization angle $\psi$ are varied while the optimal SNR is fixed at 100. The measurement errors are roughly of the same order of magnitude. However, $\sigma_{\phi}$ and $\sigma_{\psi}$ are relatively large for $\hat{\theta}=0.01\pi$, while $\sigma_{\log_{10}H}$ is relatively large for $\hat{\theta}=0.49\pi$, likely due to degeneracies in the signal at $\theta=0$ and $\theta=\pi/2$.
@@
The bottom panel of Fig. 18 shows the simulated measurement results by a single BBO detector at different SNRs. The measurement precision of $H$ and $\tau_{*}$ appears to improve with increasing SNR. Compared with the results obtained with Taiji, the agreement with the FIM estimates is better for $\rho\sim 10$, possibly due to weaker parameter degeneracies in the signal for BBO.
@@
Figs. 19-20 present example results of parameter estimation for corrected soft waveforms using an uncorrected template. In Fig. 19, we show the relative measurement errors of a soft displacement-memory signal with real linear and quadratic corrections (see Sec. IV.4), $\tilde{h}=\frac{iHe^{-2i\psi}}{2\pi f}e^{2\pi ift_{*}}[1+C(f)]$, with $C(f)=a_{0}|f|+a_{1}f^{2}$. To isolate the intrinsic parameter-estimation performance, we omit noise injection. The considered correction factors are relatively large and are plotted in Fig. 21. We find that these corrections only have a minor impact on the measurement of $\tau_{*}$, but they introduce clear biases in $\log_{10}H$. Such biases are expected, given that the high-frequency spectrum is largely modified by the correction. When the correction is moderate, however, the bias remains modest, staying within the $2\sigma$ deviation. For the other parameters, which have considerably lower measurement precision and show visible bias even in the absence of corrections, the results are not significantly affected, although the biases tend to increase.
@@
As realistic examples, we simulate the measurement of a stellar-mass compact binary scattering event with the Taiji detector and a quasi-circular stellar-mass BBH merger event with the BBO detector. For the former, we use the waveform in Fig. 4 (neglecting the real-part spectrum), and for the latter, the $q=1$ non-spinning waveform in Fig. 7. Both waveforms can be approximated by soft waveforms of displacement memory with real correction factors $C(f)$, as shown in Fig. 21. Here we choose $H=(2\pi f|\tilde{h}|)_{f=10^{-4}\,\text{Hz}}$, such that $C(10^{-4}\,\text{Hz})=0$. As above, a soft-waveform template is employed to perform the parameter estimation. The posterior distributions of the memory amplitude and polarization angle, together with the detailed parameter settings, are shown in Fig. 20. In both examples, we find that the memory parameters can be effectively constrained, with the amplitude $\log_{10}H$ determined with higher precision.